# Robust Superradiance and Spontaneous Spin Ordering in Disordered Waveguide QED

### v1

@@ IV.3 Second-Order Correlations
For later times ($t\neq 0$), we compute $g^{(2)}_{\eta,\eta^{\prime}}(t,t)$ numerically. The results are shown in Fig. 6. In the absence of disorder, both auto- and cross-correlations dip to a value of $\approx 1$ near the superradiant peak, as the light emitted is essentially coherent. Apart from a slight overall delay, the same behavior is still observed for moderate levels of disorder, despite a significant reduction of the cross-correlations at the initial stage. For strong disorder, we observe a different behavior. The dip in the autocorrelations is much less pronounced, and cross-correlations remain close to a value of $\approx 1$ during the whole decay process. The fact that the auto-correlations are consistently larger than cross-correlations implies an asymmetry in emission: a photon detected in one output channel is more likely to be followed by another photon in the same direction. This correlation has been recently studied in Ref. [28] for small periodic atomic arrays. As shown later, the persistence of this asymmetry becomes transparent and more evident when considering the homodyne trajectories.
@@
to quantify the total probability of emitting two successive photons into any direction. An initial value of $\tilde{g}^{(2)}(0,0)>1$ implies that the emission of the first photon enhances the emission probability of the second photon compared to two completely independent emission events. Consequently, the decay rate $\mathcal{R}$ initially increases as a function of time and reaches a maximum during the subsequent decay. The condition $\tilde{g}^{(2)}(0,0)>1$ has thus been proposed as a signature of superradiance [21], and more recently used to conjecture the superradiant scaling $\mathcal{R}_{\star}\sim N^{\tilde{g}^{(2)}}$ [40], which is incompatible with our numerical and analytical results.
@@ V.1 Maximal Decay Rate
$|\rho=\sum_{s}p_{s}\ket{\psi^{(s)}_{\text{prod}}}\!\bra{\psi^{(s)}_{\text{prod}}}$ (27)
@@
$\begin{split}&\frac{\gamma}{2}\braket{\psi_{\text{prod}}|J_{\rm R}^{\dagger}J_{\rm R}+J_{\rm L}^{\dagger}J_{\rm L}|\psi_{\text{prod}}}\\ &=\frac{\gamma}{4}\sum_{i,j}\cos(\xi_{i}-\xi_{j})\cos(\phi_{i}-\phi_{j})\sin\theta_{i}\sin\theta_{j}.\end{split}$ (29)
By maximizing this expression with respect to all angles $\{\theta_{j},\phi_{j}\}$, we obtain the upper bound $\mathbb{R}$, which—under the validity of Eq. (27)—represents the maximal emission rate that can possibly be reached for a given disorder configuration. Since this maximum is always reached for $\theta_{j}=\pi/2$, i.e. when the radiating dipoles point along the XY-plane of the Bloch sphere, we find
@@
Recently, by applying a theorem proved by Lieb [63] and Bravyi et al. [64] for bounding the ground state energy of spin-$1/2$ Hamiltonians, an upper bound for $\mathcal{R}_{\star}$ has been derived in Ref. [39]. By refining this calculation, and after taking the disorder-average, we obtain (see Appendix F)
@@
Instead, the bound in Eq. (30) arises from an optimization over separable states, which turns out to be a much tighter bound due to the semiclassical nature of superradiance. We first note that for weak disorder with $\Theta\leq\pi/2$, the quantities $\cos(\xi_{i}-\xi_{j})\geq 0$ are all positive. This makes $\{\phi_{i}=\phi_{j}\}$ the optimal configuration and we obtain the exact upper bound
@@
which is also very close to the numerical values for $\mathcal{R}_{\star}$ plotted in Fig. 7. Note that already in the absence of disorder, the rate $\mathcal{R}_{\star}/(\gamma N^{2})\simeq 0.19$ is slightly lower than the maximally possible rate of $\mathbb{R}/(\gamma N^{2})=1/4$, which can be attributed to the fact that the state in Eq. (27) represents a mixture of trajectories, for which the superradiant burst occurs at slightly different times [11] (see also [36] for the mixture of QSDMF trajectories). Leaving this averaging effect aside, this comparison suggests that even in the presence of disorder, the dipole tends to maximize the decay rate during the superradiant burst and that we can use $\mathbb{R}\gtrsim\mathcal{R}_{\star}$ as a tight upper bound.
@@
Under the assumption that $\mathbb{R}$ is smaller than the bound without disorder, $\gamma N^{2}/4$, we can constrain the maximal decay rate to a rather narrow interval, $\mathbb{R}_{<}\leq\mathbb{R}\leq\gamma N^{2}/4$, or $(\mathbb{R}-\mathbb{R}_{<})/(\gamma N^{2})\leq 1/8$. While this is a constraint on $\mathbb{R}$ and not on $\mathcal{R}_{\star}$, we find from numerics that the correspondence between $\mathbb{R}_{<}$ and $\mathcal{R}_{\star}$ is very close. Based on this observation we conjecture that the dipole configurations $\{|\phi_{i}-\phi_{j}|=|\xi_{i}-\xi_{j}|\}$ that enter $\mathbb{R}_{<}$ also play a dominant role in the actual dynamics.
Note that the upper bound $\mathbb{R}$ is independent of the Hamiltonian $H$, which indicates that details in $H$ do not change the superradiant scaling. For example, as shown in Appendix E, artificially ignoring the Hamiltonian only modifies the prefactor of the $N^{2}$-scaling.
Figure 7: Asymptotic limit. Plot of the scaled superradiant rate $\mathcal{R}_{\star}/(\gamma N^{2})$ for large atom numbers $N$ and as a function of the disorder strength $\theta$. Circles show the fitted values of $r_{0}$ in Eq. (18), while crosses represent the rates for a large, but finite $N=400$. The solid lines show (i) exact analytic bound derived in Eq. (32) for $\Theta<\pi/2$ and (ii) the approximate upper bound $\mathbb{R}_{<}$ given in Eq. (34). The agreement between the numerical data and $\mathbb{R}_{<}$ obtained from a product state ansatz confirms the semiclassical nature of superradiance and that the asymptotic $N^{2}$-scaling remains intact at arbitrarily large disorder strength.
@@ V.2 Spontaneous Spin Ordering
In our analysis above, we find that alignments of dipoles satisfying $\{|\phi_{i}-\phi_{j}|=|\xi_{i}-\xi_{j}|\}$ lead to close-to-maximal decay rates for a given disorder realization. Therefore, we expect that these configurations will also dominate the actual mixed state near the time of the superradiant burst. Importantly, this condition is satisfied by two sets of configurations, $\{\phi_{j}=\phi_{1}+(\xi_{j}-\xi_{1})\}$ and $\{\phi_{j}=\phi_{1}-(\xi_{j}-\xi_{1})\}$, which both fix the relative phases of the individual dipole according to their location along the waveguide [see Fig. 1(c)]. Because the system has a global $U(1)$ symmetry, the phase of the first atom, $\phi_{1}$, remains undetermined. By assuming that only those configurations contribute, we can approximate the state at the burst time $t_{\star}$ as
@@
To test the validity of the prediction for $\rho(t_{\star})$ in Eq. (35), we take a closer look at the spin orientations obtained from our QSDMF simulations. Specifically, for a given disorder realization, we evaluate all propagation phases $\Delta\xi=(\xi_{i}-\xi_{j})\bmod(2\pi)$ together with the set of relative dipole orientations, $\Delta\phi=(\phi_{i}-\phi_{j})\bmod(2\pi)$ in the QSDMF trajectories at $t=t_{\star}$. In Fig. 8, we show the resulting joint probability density $p(\Delta\xi,\Delta\phi)$ for three different strengths of the disorder. According to our analysis above, for weak disorder we expect this distribution to be peaked around $\Delta\phi=0$, which maximizes the decay rate independently of the value of $\Delta\xi$. In contrast, for a strongly disordered system, the probability density should be peaked around the values $\Delta\phi=\pm\Delta\xi$, i.e., along the diagonals of the plots.
Indeed, we find that both trends are clearly visible in the examples shown in Fig. 8, although considerably broadened by fluctuations. This confirms that, under strong disorder, the spins self-organize their orientations along a local direction determined by their random position. This is in stark contrast to the weakly disordered cases, where spins tend to synchronize with a common phase. Interestingly, under all conditions, we find a particularly strong alignment tendency for atoms that are multiples of a wavelength apart, i.e., $\Delta\xi=0$. In this case, there is an almost vanishing probability to find a nonzero relative phase $\Delta\phi$, between the corresponding dipoles.
Note that this spin ordering is not constrained to systems with disorder. In Appendix H we perform an identical analysis for periodic arrays, where for lattice spacing that do not match the emission wavelength $\lambda_{0}$, similar orderings of the atomic dipoles arise.
Figure 8: Spontaneous spin ordering. The joint probability distribution $p(\Delta\xi,\Delta\phi)$ of the relative propagating phase $\Delta\xi$ and relative dipole orientation $\Delta\phi$ are plotted for (a) $\Theta=\pi/4$, (b) $\Theta=\pi/2$, and (c) $\Theta=2\pi$. For strong disorder ($\Theta=2\pi$), the distribution aligns along the diagonals $\Delta\phi\approx\pm\Delta\xi$, consistent with the predicted spin ordering pattern. For more data on the smooth crossover from weak to strong disorders, see the Supplemental Material [65]. Simulations are performed using the QSDMF method with $N=400$ atoms and $\mathcal{N}=10^{3}$ trajectories.
Our finding of spontaneous spin ordering during superradiant decay is conceptually related to other works on synchronization in superradiance. In particular, Ref. [17] studied an array of quantum dipoles under pumping, and showed that long-range interactions and collective decay can lead to phase synchronization in the steady state. While their work focuses on synchronization to a common phase under continuous pumping in the steady states, our result reveals a distinct dynamical mechanism: during spontaneous emission without external control, the spins self-organize their orientations transiently in a disorder-dependent pattern that enhances constructive interference and preserves superradiant scaling.
@@ V.3 Mirror-Asymmetric Correlations
In Fig. 9, we plot the actual distribution of the pair of rates $(\mathcal{R}_{\text{R}},\mathcal{R}_{\text{L}})$ in each QSDMF trajectory at time $t=t_{\star}$. These plots provide a very intuitive way to illustrate the effect of mirror-asymmetric emission in superradiant decay. For weak disorder both rates are strongly correlated around $\mathcal{R}_{\text{R}}\approx\mathcal{R}_{\text{L}}$. This implies that, along each trajectory, photons are emitted approximately equally likely to the left and to the right. In the regime of strong disorder, the rates are anti-correlated with $\mathcal{R}_{\text{R}}+\mathcal{R}_{\text{L}}\approx\mathcal{R}_{\star}$. Therefore, in this regime, the atomic dipoles evolve into configurations that successively lead to preferred emission into either the right or the left. Note that, instead of keeping a preferred emission direction in time evolution, the atomic dipoles can change their preferred emission direction during the decay in the individual trajectories, as shown in Appendix G.
In summary, Fig. 8 and Fig. 9 clearly highlight the difference in the spontaneously established ordering of the atomic dipoles in weakly and strongly disordered arrays. For strong disorder, the peaks in the diagonals of the distribution of $p(\Delta\xi,\Delta\phi)$ in Fig. 8(c) and the anticorrelations seen in the rates in Fig. 9(b) also demonstrate the relevance of those dipole orderings that maximize emission to the left or the right. However, as evident from the photon correlations and the distribution of spin orientations, these orderings are distributions with a considerable broadening around the optimal configurations, which is also the reason underlying the discrepancy between the upper bound $\mathbb{R}_{<}$ and actual numerical data shown in Fig. 7.
@@ VI.1 Inhomogeneous Broadening
Indeed, in Fig. 13, we confirm this argument by simulating the effect of frequency disorder of increasing strength using the DTWA method. While there is a visible influence of inhomogeneous broadening when $N$ is moderately large, the deviations from the case without disorder vanish for sufficiently large $N$ even for very strong spectral disorder (see more analysis in the Supplemental Material [65]).
@@ VI.2 Non-Markovian Effects
Apart from various types of spatial and spectral disorders, another relevant question about superradiant decay concerns the role of non-Markovian effects. The non-Markovian regime has been notoriously difficult to study even for a few emitters [66, 67], let alone large-scale simulations of $N$ emitters. Non-Markovian superradiance has only been explored in very special cases, such as the cascaded system with time delays [34], where an effective master equation can be derived [68, 69]. In the current setting, non-Markovianity can be explored in terms of the extended master equation introduced Eq. (7), where the collective decay channels are implemented via coherent coupling to two lossy cavity modes. In this setting, Markovianity and thus a recovery of the original master equation in Eq. (1) is ensured when the dynamics of the atoms are slow compared to the decay times of the cavities, $\kappa^{-1}$. During the initial stage of the decay, this condition is satisfied when $\kappa\gtrsim\gamma N$. However, the situation is less clear around the time $t_{\star}$ of the superradiant burst, where, in view of the scaling $\mathcal{R}_{\star}\sim\gamma N^{2}$, the accelerated atomic decay may eventually be faster than the cavity loss. Specifically, under the condition $\gamma N\ll\kappa\ll\gamma N^{2}$, which can be encountered for sufficiently large $N$, the superradiant dynamics may both be slow and fast compared to the photonic environment, and the validity of a Markovian description is a priori unclear.
@@
These numerical results demonstrate that superradiance is robust against weak non-Markovianity, and clarify that the enhanced rate $\mathcal{R}_{\star}$ does not affect the validity of a Markovian description for superradiant decay as long as $\kappa\gg\gamma N$. As discussed in Appendix B, we can reconcile this finding by noticing that the conditions for Markovianity from above are unnecessarily stringent, since the photons can leave the cavity modes at a rate $\sim\kappa N$ with $\mathcal{O}(N)$ photons inside. The relevant time scales are set by the atomic dynamics, which evolve on a timescale of order $1/\gamma$ initially and $1/(\gamma N)$ during superradiance. Consequently, it suffices to require $\kappa\gg\gamma$ initially and $\kappa\gg\gamma N$ in the superradiant regime to maintain Markovianity. Hence, the ratio $\gamma N/\kappa$ serves as a natural control parameter for the degree of non-Markovianity in the system.
@@
Figure 11: Impact of non-Markovianity on superradiant scaling. (a) Benchmark of the DTWA method against exact diagonalization (ED) of $N=5,10,15$ atoms coupled to one lossy cavity. (b) Scaling of the superradiant rate $\mathcal{R}_{\star}/(\gamma N^{2})$ for $\kappa/(\gamma N)=1,2,5,10,50$ (from bottom to top). The inset shows the value of $\mathcal{R}_{\star}/(\gamma N^{2})$ for $N=400$ as a function of $\gamma N/\kappa$. Even in the non-Markovian regime [small values of $\kappa/(\gamma N)$], the $N^{2}$-scaling persists. The prefactor of $N^{2}$ scaling increases to the factor $\approx 0.2$ of the Dicke case as $\kappa$ increases. All results have been obtained by averaging over $10^{3}$ trajectories.
VII Conclusions

### v2

@@ IV.3 Second-Order Correlations
For later times ($t\neq 0$), we compute $g^{(2)}_{\eta,\eta^{\prime}}(t,t)$ numerically. The results are shown in Fig. 6. In the absence of disorder, both auto- and cross-correlations dip to a value of $\approx 1$ near the superradiant peak, as the light emitted is essentially coherent. Apart from a slight overall delay, the same behavior is still observed for moderate levels of disorder, despite a significant reduction of the cross-correlations at the initial stage. For strong disorder, we observe a different behavior. The dip in the autocorrelations is much less pronounced, and cross-correlations remain close to a value of $\approx 1$ during the whole decay process. The fact that the auto-correlations are consistently larger than cross-correlations implies an asymmetry in emission: a photon detected in one output channel is more likely to be followed by another photon in the same direction. This correlation has been recently studied in Ref. [28] for small periodic atomic arrays. As shown later, the persistence of this asymmetry becomes transparent and more evident when considering the heterodyne trajectories.
@@
to quantify the total probability of emitting two successive photons into any direction. An initial value of $\tilde{g}^{(2)}(0,0)>1$ implies that the emission of the first photon enhances the emission probability of the second photon compared to two completely independent emission events. Consequently, the decay rate $\mathcal{R}$ initially increases as a function of time and reaches a maximum during the subsequent decay. The condition $\tilde{g}^{(2)}(0,0)>1$ has thus been proposed as a signature of superradiance [21], and more recently used to conjecture the superradiant scaling $\mathcal{R}_{\star}\sim N^{\tilde{g}^{(2)}}$ [40].
@@ V.1 Maximal Decay Rate
Figure 7: Asymptotic limit. Plot of the scaled superradiant rate $\mathcal{R}_{\star}/(\gamma N^{2})$ for large atom numbers $N$ and as a function of the disorder strength $\Theta$. Circles show the fitted values of $r_{0}$ in Eq. (18), while crosses represent the rates for a large, but finite $N=400$. The solid lines show (i) the exact analytic bound derived in Eq. (32) for $\Theta<\pi/2$ and (ii) the approximate upper bound $\mathbb{R}_{<}$ given in Eq. (34). The qualitative agreement between the numerical data and $\mathbb{R}_{<}$ obtained from a product state ansatz confirms the semiclassical nature of superradiance and that the asymptotic $N^{2}$-scaling remains intact at arbitrarily large disorder strength.
@@
$|.\rho=\sum_{s}p_{s}\ket{\psi^{(s)}_{\text{prod}}}\!\bra{\psi^{(s)}_{\text{prod}}}$. (27)
@@
$\begin{split}&\frac{\gamma}{2}\braket{\psi_{\text{prod}}|\left(J_{\rm R}^{\dagger}J_{\rm R}+J_{\rm L}^{\dagger}J_{\rm L}\right)|\psi_{\text{prod}}}\\ &=\frac{\gamma}{4}\sum_{i,j}\cos(\xi_{i}-\xi_{j})\cos(\phi_{i}-\phi_{j})\sin\theta_{i}\sin\theta_{j}.\end{split}$ (29)
By maximizing this expression with respect to all angles $\{\theta_{j},\phi_{j}\}$, we obtain the upper bound $\mathbb{R}$, which—under the validity of Eq. (27)—represents the maximal emission rate that can possibly be reached for a given disorder configuration. Since this maximum is always reached for $\theta_{j}=\pi/2$, i.e., when the radiating dipoles point along the XY-plane of the Bloch sphere, we find
@@
Recently, by applying a theorem proved by Lieb [67] and Bravyi et al. [68] for bounding the ground state energy of spin-$1/2$ Hamiltonians, an upper bound for $\mathcal{R}_{\star}$ has been derived in Ref. [39]. By refining this calculation, and after taking the disorder average, we obtain (see Appendix E)
@@
Instead, the bound in Eq. (30) arises from an optimization over separable states, which turns out to be a much tighter bound. We first note that for weak disorder with $\Theta\leq\pi/2$, the quantities $\cos(\xi_{i}-\xi_{j})\geq 0$ are all positive. This makes $\{\phi_{i}=\phi_{j}\}$ the optimal configuration and we obtain the exact upper bound
@@
which is also very close to the numerical values for $\mathcal{R}_{\star}$ plotted in Fig. 7. Note that already in the absence of disorder, the rate $\mathcal{R}_{\star}/(\gamma N^{2})\simeq 0.19$ is slightly lower than the maximally possible rate of $\mathbb{R}/(\gamma N^{2})=1/4$, which can be attributed to the fact that the state in Eq. (27) represents a mixture of trajectories, for which the superradiant burst occurs at slightly different times [11] (see also [36] for the mixture of QSDMF trajectories). Leaving this averaging effect aside, this comparison suggests that even in the presence of disorder, the dipole tends to maximize the decay rate during the superradiant burst and that we can use $\mathbb{R}\gtrsim\mathcal{R}_{\star}$ as a rather tight upper bound.
@@
Under the assumption that $\mathbb{R}$ is smaller than the bound without disorder, $\gamma N^{2}/4$, we can constrain the maximal decay rate to a rather narrow interval, $\mathbb{R}_{<}\leq\mathbb{R}\leq\gamma N^{2}/4$, or $(\mathbb{R}-\mathbb{R}_{<})/(\gamma N^{2})\leq 1/8$. While this is a constraint on $\mathbb{R}$ and not on $\mathcal{R}_{\star}$, we find from numerics that the correspondence between $\mathbb{R}_{<}$ and $\mathcal{R}_{\star}$ is very close. The discrepancy can be attributed to the distribution around the optimal configurations in both the azimuthal angles $\phi_{j}$ and the polar angles $\theta_{j}$ (see the Supplemental Material [60]), which is neglected in deriving the upper bound. Based on this observation we conjecture that the dipole configurations $\{|\phi_{i}-\phi_{j}|=|\xi_{i}-\xi_{j}|\}$ that enter $\mathbb{R}_{<}$ also play a dominant role in the actual dynamics. This conjecture is also supported by the basic observation that $\mathcal{R}_{\star}$ saturates above a value of $\Theta\gtrsim\pi$. At this value, the distribution of twice the phase differences, $2(\xi_{i}-\xi_{j})\bmod(2\pi)$, which enters the approximated expression in Eq. (33), becomes flat. Instead, for a decay rate that only depends on the actual propagation phases, $\Delta\xi_{ij}=(\xi_{i}-\xi_{j})\bmod(2\pi)$, a saturation of $\mathcal{R}_{\star}$ would be expected for $\Theta\gtrsim 2\pi$ (see the Supplemental Material [60]).
Note that the upper bound $\mathbb{R}$ is independent of the Hamiltonian $H$, which indicates that details in $H$ do not change the superradiant scaling. For example, as shown in Appendix F, artificially ignoring the Hamiltonian only modifies the prefactor of the $N^{2}$-scaling.
@@ V.2 Spontaneous Spin Ordering
Figure 8: Spontaneous spin ordering. The joint probability distribution $p(\Delta\xi,\Delta\phi)$ of the relative propagating phase $\Delta\xi$ and relative dipole orientation $\Delta\phi$ are plotted for (a) $\Theta=\pi/4$, (b) $\Theta=\pi/2$, and (c) $\Theta=2\pi$. For strong disorder ($\Theta=2\pi$), the distribution aligns along the diagonals $\Delta\phi\approx\pm\Delta\xi$, consistent with the predicted spin ordering pattern. For more data on the crossover between weakly and strongly disordered configurations, see the Supplemental Material [60]. Simulations are performed using the QSDMF method with $N=400$ atoms and $\mathcal{N}=10^{3}$ trajectories.
In our analysis above, we find that alignments of dipoles satisfying $\{|\phi_{i}-\phi_{j}|=|\xi_{i}-\xi_{j}|\}$ lead to close-to-maximal decay rates for a given disorder realization. Therefore, we expect that these configurations will also dominate the actual mixed state near the time of the superradiant burst. Importantly, this condition is satisfied by two sets of configurations, $\{\phi_{j}=\phi_{1}+(\xi_{j}-\xi_{1})\}$ and $\{\phi_{j}=\phi_{1}-(\xi_{j}-\xi_{1})\}$, which both fix the relative phases of the individual dipole according to their location along the waveguide [see Fig. 1(b)]. Because the system has a global $U(1)$ symmetry, the phase of the first atom, $\phi_{1}$, remains undetermined. By assuming that only those configurations contribute, we can approximate the state at the burst time $t_{\star}$ as
@@
To test the validity of the prediction for $\rho(t_{\star})$ in Eq. (35), we take a closer look at the spin orientations obtained from our QSDMF simulations. Specifically, for a given disorder realization, we evaluate all propagation phases $\Delta\xi_{ij}=(\xi_{i}-\xi_{j})\bmod(2\pi)$ together with the set of relative dipole orientations, $\Delta\phi_{ij}=(\phi_{i}-\phi_{j})\bmod(2\pi)$ in the QSDMF trajectories at $t=t_{\star}$. In Fig. 8, we show the resulting joint probability density $p(\Delta\xi,\Delta\phi)$ for three different strengths of the disorder. According to our analysis above, for weak disorder we expect this distribution to be peaked around $\Delta\phi=0$, which maximizes the decay rate independently of the value of $\Delta\xi$. In contrast, for a strongly disordered system, the probability density should be peaked around the values $\Delta\phi=\pm\Delta\xi$, i.e., along the diagonals of the plots.
Indeed, we find that both trends are clearly visible in the examples shown in Fig. 8, although considerably broadened by fluctuations. This confirms that, under strong disorder, the spins self-organize their orientations along a local direction determined by their random position. This is in stark contrast to the weakly disordered cases, where spins tend to synchronize with a common phase. Interestingly, under all conditions, we find a particularly strong alignment tendency for atoms that are multiples of a wavelength apart, i.e., $\Delta\xi=0$. In this case, there is an almost vanishing probability to find a nonzero relative phase $\Delta\phi$, between the corresponding dipoles. Note that these findings are not constrained to systems with disorder. In Appendix G, we perform an identical analysis for periodic arrays, where for lattice spacing that do not match the emission wavelength $\lambda_{0}$, similar orderings of the atomic dipoles arise.
Our finding of spontaneous spin ordering during superradiant decay is conceptually related to other works on synchronization in superradiance. In particular, Ref. [17] studied an array of quantum dipoles under continuous pumping, and showed that long-range interactions and collective decay can lead to synchronization to a common phase in the steady state. Instead, our result reveals a distinct dynamical mechanism: during spontaneous emission without external control, the spins self-organize their orientations transiently in a disorder-dependent pattern that enhances constructive interference and preserves superradiant scaling.
@@ V.3 Mirror-Asymmetric Correlations
In Fig. 9, we plot the actual distribution of the pair of rates $(\mathcal{R}_{\text{R}},\mathcal{R}_{\text{L}})$ in each QSDMF trajectory at time $t=t_{\star}$. These plots provide a very intuitive way to illustrate the effect of mirror-asymmetric emission in superradiant decay. For weak disorder both rates are strongly correlated around $\mathcal{R}_{\text{R}}\approx\mathcal{R}_{\text{L}}$. This implies that, along each trajectory, photons are emitted approximately equally likely to the left and to the right. In the regime of strong disorder, the rates are anti-correlated with $\mathcal{R}_{\text{R}}+\mathcal{R}_{\text{L}}\approx\mathcal{R}_{\star}$. Therefore, in this regime, the atomic dipoles evolve into configurations that successively lead to preferred emission into either the right or the left. Additional simulations presented in Appendix H show that this preference exists during most parts of the decay, although individual trajectories may also change their preferred emission direction during the process.
In summary, Fig. 8 and Fig. 9 clearly highlight the difference in the spontaneously established ordering of the atomic dipoles in weakly and strongly disordered arrays. For strong disorder, the peaks in the diagonals of the distribution of $p(\Delta\xi,\Delta\phi)$ in Fig. 8(c) and the anticorrelations seen in the rates in Fig. 9(b) also demonstrate the relevance of those dipole orderings that maximize emission to the left or the right. However, as evident from the photon correlations and the distribution of spin orientations, these orderings are distributions with a considerable broadening around the optimal configurations, which is also the reason underlying the discrepancy between the upper bound $\mathbb{R}_{<}$ and actual numerical data shown in Fig. 7. Note that our analysis has focused on spin ordering in the bulk of the atomic array. In a related recent analysis of spin-ordering in pumped atomic arrays [69], it was observed that atoms at the left (right) end align anticlockwise (clockwise) and therefore prefer emission to the left (right). In our setup, such boundary effects also exist, but their contribution to the overall dynamics and decay asymmetry is minor (see Appendix I).
@@ VI.1 Inhomogeneous Broadening
Indeed, in Fig. 10, we confirm this argument by simulating the effect of frequency disorder of increasing strength using the DTWA method. While there is a visible influence of inhomogeneous broadening when $N$ is moderately large, the deviations from the case without disorder vanish for sufficiently large $N$ even for very strong spectral disorder (see more detailed analysis in the Supplemental Material [60]).
@@ VI.2 Non-Markovian Effects
Apart from various types of spatial and spectral disorders, another relevant question about superradiant decay concerns the role of non-Markovian effects. The non-Markovian regime has been notoriously difficult to study even for a few emitters [70, 71], let alone large-scale simulations of $N$ emitters. Non-Markovian superradiance has only been investigated in very special cases, such as the cascaded system with time delays [34], where an effective master equation can be derived [72, 73]. In the current setting, non-Markovianity can be explored in terms of the extended master equation introduced Eq. (7), where the collective decay channels are implemented via coherent coupling to two lossy cavity modes. In this setting, Markovianity and thus a recovery of the original master equation in Eq. (1) is ensured when the dynamics of the atoms are slow compared to the decay times of the cavities, $\kappa^{-1}$. During the initial stage of the decay, this condition is satisfied when $\kappa\gtrsim\gamma N$. However, the situation is less clear around the time $t_{\star}$ of the superradiant burst, where, in view of the scaling $\mathcal{R}_{\star}\sim\gamma N^{2}$, the accelerated atomic decay may eventually be faster than the cavity loss. Specifically, under the condition $\gamma N\ll\kappa\ll\gamma N^{2}$, which can be encountered for sufficiently large $N$, the superradiant dynamics may both be slow and fast compared to the photonic environment, and the validity of a Markovian description is a priori unclear.
@@
These numerical results demonstrate that superradiance is robust against weak non-Markovianity, and clarify that the enhanced rate $\mathcal{R}_{\star}$ does not affect the validity of a Markovian description for superradiant decay as long as $\kappa\gg\gamma N$. We can reconcile this finding by noticing that the conditions for Markovianity from above are unnecessarily stringent, since the photons can leave the cavity modes at a rate $\sim\kappa N$ with $\mathcal{O}(N)$ photons inside. The relevant time scales are set by the atomic dynamics, which evolve on a timescale of order $1/\gamma$ initially and $1/(\gamma N)$ during superradiance. Consequently, it suffices to require $\kappa\gg\gamma$ initially and $\kappa\gg\gamma N$ in the superradiant regime to maintain Markovianity. Hence, the ratio $\gamma N/\kappa$ serves as a natural control parameter for the degree of non-Markovianity in the system.
@@
Figure 11: Impact of non-Markovianity on superradiant scaling. (a) Benchmark of the DTWA method against exact diagonalization (ED) of $N=5,10,15$ atoms coupled to one lossy cavity. (b) Scaling of the superradiant rate $\mathcal{R}_{\star}/(\gamma N^{2})$ for $\kappa/(\gamma N)=1,2,5,10,50$ (from bottom to top). The inset shows the value of $\mathcal{R}_{\star}/(\gamma N^{2})$ for $N=400$ as a function of $\gamma N/\kappa$. Even in the non-Markovian regime [small values of $\kappa/(\gamma N)$], the $N^{2}$-scaling persists. The prefactor of $N^{2}$ scaling increases to the factor $\approx 0.2$ of the Dicke case as $\kappa$ increases. All results have been obtained by averaging over $\mathcal{N}=10^{3}$ trajectories.
VII Conclusion and Outlook